# Through the Magnifying Glass: Adaptive Perception Magnification for Hallucination-Free VLM Decoding

### v2

@@ 5 Conclusion
We introduce the Perception Magnifier (PM), a visual decoding method that refines image inputs to mitigate hallucinations in vision-language models. PM adaptively enlarges visually critical regions identified by multi-layer attention iteratively, providing enhanced effective resolution on relevant finer-grained details at each decoding step. Experiments on MME, POPE, and LLaVA-Bench demonstrate that PM significantly reduces hallucinations at QA and open-end generation, while preserving reasoning capabilities. Future work may focus on optimizing the decision process for input magnification at each step, refining the identification of salient tokens, and extending the magnification-based approach to a broader range of VLM architectures.
Supplementary Material
@@
6 Datasets
MME [20] benchmark is designed to comprehensively assess both perception and cognition abilities of VLMs across 14 distinct subtasks on 2,370 questions. In our study, we focus primarily on the Perception subset of MME following [34] to evaluate the effectiveness of our method in controlling hallucinations.
POPE (Polling-based Object Probing Evaluation) [37] is a 9 $\times$ 3,000 question benchmark for evaluating object hallucination in VLMs. It formulates hallucination evaluation as a binary classification task by posing yes/no questions about object presence, addressing limitations of CHAIR [52], which relies on exact text matching, and providing a more robust evaluation less sensitive to prompt variations.
LlaVA-Bench [40] is a benchmark designed to evaluate VLMs in real-world visual chat scenarios. It comprises two components: LLaVA-Bench (COCO), utilizing images from the COCO dataset, and LLaVA-Bench (In-the-Wild), a more challenging subset featuring 24 diverse images—such as indoor and outdoor scenes, memes, paintings, and sketches—paired with 60 detailed, manually curated questions. The benchmark covers various vision-language tasks, making it suitable for assessing model generalization. In our study, we employ both COCO and In-the-Wild splits to comprehensively evaluate different methods on hallucination control, cognitive abilities, and performance in open-ended question answering.
7 Baselines
OPERA [24] is a beam-search-based method, examining the presence of a hallucinate beam and pruning the beam based on its attention on information aggregation patterns.
VCD [34] introduces noise into the input image and processes it through the VLM to generate a hallucinative output. By leveraging contrastive decoding (CD) to penalize high-probability predictions that align with the hallucinative output, VCD reduces the likelihood of hallucinations in the final prediction, resulting in more reliable responses.
ICD [62] generates a hallucinative output for CD-based decoding by appending ”You are a confused object detector” to the input prompt.
SID [25] induces a hallucinative output for CD-based decoding by identifying key visual tokens in the image and pruning them through forward.
VDD [79] mitigates language-induced hallucinations in VLMs by generating a hallucinative output through the removal or replacement of visual tokens with <unk> tokens.
PAI [43] enhances visual faithfulness by increasing the weights of visual tokens through attention mechanisms while penalizing outputs generated without visual inputs using CD.
IBD [84] also biases the output toward visual inputs by adding visual attentions and refines it against the original output using CD. Additionally, it adaptively adjusts the contrastive update strength at each decoding step based on contextual needs.
API [73] highlights key visual regions by overlaying a gray-level mask on the input image. The mask is guided by a heatmap derived from image-text associations using a CLIP model (-C) or a LLaVA model (-L).
ViCrop [76] estimates a bounding box around the key region in the image and generates a new input by cropping and resizing this region. Three approaches are proposed for bounding box estimation with different level of reliance on gradient information: relative attention (-R), gradient attention (-G), and pure gradient (-P). The pure-gradient approach is not benchmarked due to its higher GPU memory consumption, which necessitates lower model precision and results in degraded performance, making it unsuitable for a fair comparison.
8 Performance on Shikra
To assess the generalizability of PM beyond LLaVA, we apply it, along with several comparative methods, to Shikra-7B [6], a VLM designed for referential dialogue. The performance is evaluated on MME and POPE using Shikra with iterative refinement threshold $\beta=1.5$. The results are presented in Tab. 7. Note that certain comparative methods are excluded either due to the lack of available implementations for Shikra or constraints on computational resources.
@@ 9 Qualitative Samples on PM
Figure 4: Additional qualitative examples on the POPE benchmark showing the input question, the original image, the processed perception map $\mathcal{H}_{\text{proc}}$, and the perception-magnified image, arranged from left to right. The answers generated by greedy decoding and PM are compared, with correct answers highlighted in green and incorrect ones in red.
Fig. 4 presents additional qualitative examples illustrating the input question, the original image, the processed perception map $\mathcal{H}_{\text{proc}}$, and the perception-magnified image. The results demonstrate that the proposed method effectively enlarges small objects relevant to the question while preserving fine details. In many cases, the VLM benefits from the perception-magnified image, as it can now recognize previously overlooked objects that were too small to be detected in the original image. This leads to more accurate predictions, as observed in the first and second rows.
@@ 10 Limitations
Although region magnification allows the model to observe fine-grained visual details without altering the overall image structure, it may distort local shape properties. Specifically, in tasks requiring precise shape or length recognition, such as distinguishing letters or assessing object sizes, PM does not guarantee proportional magnification. This can introduce challenges for accurate recognition and induces hallucinations. For instance, when applied to CLEVR-style questions, as shown in Fig. 5, with the query “Please estimate how many times large is the blue sphere compared to the red cube?”, PM magnifies both objects, causing the blue sphere to appear visually larger while distorting the shape of the other two cubes. Such unintended modifications may negatively impact final predictions. Future work could explore refinements in perception map construction or post-processing techniques to achieve more structure-preserving magnification.
@@
While the proposed method can be easily applied to interleaved VLMs with visual tokens corresponding to all patch features, applying it to VLMs where the correlation between visual tokens and image regions are non-trivial, can be slightly more complicated. The method depends on certain mapping from the visual token attention weights to the attention span on the original image. Therefore, when apply PM to models where the visual tokens are constructed with not only a feature extractor, but also a token sampler or cross-attention feature extractors, PM will need some attention mapping mechanism between visual tokens and patches. However, we can resolve this problem by introducing gradient backpropagation to address the implicit relationships between the image and tokens, refining the visual attentions and construct the perception map, similar to method proposed in [4]. Overall, such limitation does not reduce the contribution of this work, as the idea of adaptive-magnification can also always be used in aligned with interpretability tools designed for different VLMs. Alternatively, one could use PM in ways similar to [73], where any compatible model could be used as external guidance to the decoding-adaptive magnification.
11 Prompts for GPT-Assisted Evaluation
We adopt the GPT-assisted evaluation prompts directly from prior studies. For reference-based evaluation, we follow the prompt used in [40], with an example of the prompt and corresponding evaluation result provided in Tab. 8. For vision-based evaluation, we use the prompt from [24], with an example shown in Tab. 9.
Prompt for Reference-based Evaluation [Context] An aerial view of Diamond Head in the Hawaiian Islands. [Question] What is the name of this famous sight in the photo? [Assistant 1] The famous sight in the photo is Diamond Head. [End of Assistant 1] [Assistant 2] The famous sight in the photo is the Na Pali Coast, which is a scenic coastal area on the island of Kaua’i in Hawaii. [End of Assistant 2] [System] We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above. The user asks the question on observing an image. For your reference, the visual content in the image is represented with a few sentences describing the image. Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance. Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgement. Response: 10 2 Assistant 1 correctly identified the famous sight in the photo as Diamond Head, which is consistent with the context provided. The response is accurate, relevant, and directly answers the question with the appropriate level of detail given the context. Assistant 2, on the other hand, incorrectly identified the sight as the Na Pali Coast, which is not mentioned in the context. This response is inaccurate and irrelevant to the question, as it does not align with the description of the image provided. Additionally, the details about the Na Pali Coast, while informative, are not applicable to the context of Diamond Head. Therefore, Assistant 2’s response is not helpful in this scenario.
@@ 11 Prompts for GPT-Assisted Evaluation
Prompt for Vision-based Evaluation [Uncaptioned image] You are required to score the performance of two AI assistants in describing a given image. You should pay extra attention to the hallucination, which refers to the part of descriptions that are inconsistent with the image content, such as claiming the existence of something not present in the image or describing incorrectly in terms of the counts, positions, or colors of objects in the image. Please rate the responses of the assistants on a scale of 1 to 10, where a higher score indicates better performance, according to the following criteria: 1: Accuracy: whether the response is accurate with respect to the image content. Responses with fewer hallucinations should be given higher scores. 2: Detailedness: whether the response is rich in necessary details. Note that hallucinated descriptions should not count as necessary details. Please output the scores for each criterion, containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. Following the scores, please provide an explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgement. [Assistant 1] The intended effect of this painting might be to create a humorous and whimsical twist on the famous Mona Lisa artwork by featuring a dog dressed in her attire, which can evoke feelings of amusement and curiosity. [End of Assistant 1] [Assistant 2] The intended effect of this painting is to create a humorous and whimsical scene by combining a dog with a human-like appearance, such as a Renaissance-style painting. The dog is wearing a Renaissance-style dress and a hat, which adds a touch of humor and creativity to the image. The painting might be intended to evoke amusement and curiosity in the viewer, as it is not common to see a dog dressed in such a manner. The artist might have wanted to challenge the viewer’s expectations and create a unique, memorable piece of art. [End of Assistant 2] Output format: Accuracy: <Scores of the two answers> Reason: Detailedness: <Scores of the two answers> Reason: Response: ”9”: ”**Accuracy: 8 9** *Reason:** Assistant 1 accurately describes the image as a humorous twist on the Mona Lisa featuring a dog, but lacks specific details about the attire. Assistant 2 provides a more detailed and accurate description, mentioning the Renaissance-style dress and the humorous intent, which aligns well with the image content. *Detailedness: 6 9** *Reason:** Assistant 1 gives a brief description focusing on the humorous aspect but lacks detail about the attire and style. Assistant 2 offers a richer description, noting the Renaissance-style dress and hat, and elaborates on the artistic intent, providing a more comprehensive understanding of the image.”

### v3

@@ Conclusion
We introduce the Perception Magnifier (PM), a visual decoding method that refines image inputs to mitigate hallucinations in vision-language models. PM adaptively enlarges visually critical regions identified by multi-layer attention iteratively, providing enhanced effective resolution on relevant finer-grained details at each decoding step. Experiments demonstrate that PM significantly reduces hallucinations at VQA and open-end generation, while preserving reasoning capabilities. Future work may focus on optimizing the efficiency for PM at long-text generation, and mitigate shape distortions in the magnification process.
Supplementary Materials
Method AOKVQA GQA COCO
Adversarial Popular Random Adversarial Popular Random Adversarial Popular Random
Greedy 78.67 84.93 88.70 80.40 83.27 89.50 83.50 85.57 86.80
Opera 78.60 84.93 88.73 80.43 83.33 89.53 83.47 85.57 86.83
VCD 79.10 84.93 89.37 79.63 81.70 88.07 83.53 85.63 87.67
ICD 79.07 85.40 88.23 80.97 84.77 88.93 83.20 85.23 86.33
SID 78.90 85.40 88.87 80.60 83.33 88.57 83.90 86.07 87.47
VDD 81.03 88.27 90.90 82.00 85.27 89.30 85.00 87.00 88.13
M3ID 80.87 87.97 90.73 81.67 84.93 89.77 85.10 86.93 88.17
PAI 79.63 86.27 90.07 79.90 82.30 88.10 85.03 86.97 87.97
IBD 80.60 87.47 91.03 81.30 84.10 89.00 84.90 86.97 88.07
API-C 81.00 88.20 91.40 81.33 84.27 89.53 85.47 87.53 88.93
API-L 80.80 87.70 91.07 81.00 83.57 88.33 84.53 86.97 88.30
ViCrop-R 80.00 86.37 90.23 81.60 83.83 90.13 85.60 87.70 89.00
ViCrop-G 80.97 86.77 90.53 81.67 83.93 89.93 85.13 87.70 89.10
[…]
Table 5: Evaluation results for POPE accuracy with LLaVA 1.5.
Datasets
MME (Fu et al. 2023) benchmark is designed to comprehensively assess both perception and cognition abilities of VLMs across 14 distinct subtasks on 2,370 questions. In our study, we focus primarily on the Perception subset of MME following (Leng et al. 2024) to evaluate the effectiveness of our method in controlling hallucinations.
POPE (Polling-based Object Probing Evaluation) (Li et al. 2023c) is a 9 $\times$ 3,000 question benchmark for evaluating object hallucination in VLMs. It formulates hallucination evaluation as a binary classification task by posing yes/no questions about object presence, addressing limitations of CHAIR (Rohrbach et al. 2018), which relies on exact text matching, and providing a more robust evaluation less sensitive to prompt variations.
LlaVA-Bench (Liu et al. 2023) is a benchmark designed to evaluate VLMs in real-world visual chat scenarios. It comprises two components: LLaVA-Bench (COCO), utilizing images from the COCO dataset, and LLaVA-Bench (In-the-Wild), a more challenging subset featuring 24 diverse images—such as indoor and outdoor scenes, memes, paintings, and sketches—paired with 60 detailed, manually curated questions. The benchmark covers various vision-language tasks, making it suitable for assessing model generalization. In our study, we employ both COCO and In-the-Wild splits to comprehensively evaluate different methods on hallucination control, cognitive abilities, and performance in open-ended question answering.
Baselines
OPERA (Huang et al. 2024) is a beam-search-based method, examining the presence of a hallucinate beam and pruning the beam based on its attention on information aggregation patterns.
VCD (Leng et al. 2024) introduces noise into the input image and processes it through the VLM to generate a hallucinative output. By leveraging contrastive decoding (CD) to penalize high-probability predictions that align with the hallucinative output, VCD reduces the likelihood of hallucinations in the final prediction, resulting in more reliable responses.
ICD (Wang et al. 2024a) generates a hallucinative output for CD-based decoding by appending “You are a confused object detector” to the input prompt.
SID (Huo et al. 2024) induces a hallucinative output for CD-based decoding by identifying key visual tokens in the image and pruning them through forward.
VDD (Zhang et al. 2024) mitigates language-induced hallucinations in VLMs by generating a hallucinative output through the removal or replacement of visual tokens with <unk> tokens.
M3ID (Favero et al. 2024) is similar to VDD but with weighted strength of incorporating contrastive corrections based on the text-image mutual information.
PAI (Liu, Zheng, and Chen 2025) enhances visual faithfulness by increasing the weights of visual tokens through attention mechanisms while penalizing outputs generated without visual inputs using CD.
IBD (Zhu et al. 2024) also biases the output toward visual inputs by adding visual attentions and refines it against the original output using CD. Additionally, it adaptively adjusts the contrastive update strength at each decoding step based on contextual needs.
API (Yu, Yu, and Wang 2024) highlights key visual regions by overlaying a gray-level mask on the input image. The mask is guided by a heatmap derived from image-text associations using a CLIP model (-C) or a LLaVA model (-L).
ViCrop (Zhang et al. 2025a) estimates a bounding box around the key region in the image and generates a new input by cropping and resizing this region. Three approaches are proposed for bounding box estimation with different level of reliance on gradient information: relative attention (-R), gradient attention (-G), and pure gradient (-P). The pure-gradient approach is not benchmarked due to its higher GPU memory consumption, which necessitates lower model precision and results in degraded performance, making it unsuitable for a fair comparison.
Computing Infrastructure
All experiments were conducted on a single machine equipped with an NVIDIA GeForce RTX 3090 GPU and an Intel(R) Core(TM) i7-10700 CPU. The system runs Ubuntu 22.04 LTS and uses CUDA version 12.1 for GPU acceleration. Experiments were implemented in Python using PyTorch (v2.4.0) and Transformers (v4.37.2).
Maintaining Cognition Ability
MME-Cognition CS NR Trans CR Total*
Greedy 112.86 70.00 107.50 67.50 357.86
OPERA 112.86 47.50 90.00 60.00 310.36
VCD 120.00 37.50 97.50 72.50 327.50
ICD 120.71 70.00 92.50 65.00 348.21
SID 122.86 55.00 85.00 65.00 327.86
VDD 110.14 58.00 68.00 86.00 322.64
M3ID 110.86 61.00 92.50 75.00 339.36
PAI 116.43 70.00 100.00 67.50 353.93
IBD 117.14 70.00 107.50 57.50 352.14
API-C 115.00 47.50 107.50 47.50 317.50
API-L 120.00 45.00 115.00 45.00 325.00
ViCrop-R 115.00 50.00 105.00 65.00 335.00
ViCrop-G 109.29 50.00 105.00 86.00 350.29
[…]
To investigate this, we evaluate our method and baselines on the Cognition split of MME, where answering requires reasoning beyond direct visual cues. As shown in Tab. 6, most CD-based methods that explicitly modify output logits exhibit a notable decline in cognition ability, aligning with expectations. Surprisingly, PAI and IBD, which enhance the attention weight of visual tokens through intermediate embedding modifications, do not incur a substantial loss in cognition performance. This suggests that the bias introduced by their layer-wise manipulation is relatively mild and can be compensated for by subsequent model layers. Furthermore, despite API-based methods modify the image without altering the VLM’s internal reasoning, they still lead to degraded performance. We hypothesize that the gray masking applied to images in the cognition set where backgrounds are typically uniform may introduce additional perceptual burden, ultimately impacting overall performance.
Quantitative Results on POPE
Table 5 presents the detailed evaluation results across each split of the POPE benchmark. Overall, the proposed PM consistently outperforms existing methods on all splits, demonstrating its effectiveness in mitigating hallucination in yes/no existence-based VQA. While the improvement on the ‘popular’ split of GQA is relatively smaller, this may be attributed to the inherently dense scenes and lower object saliency in the GQA benchmark, which can pose challenges for our magnification-based visual enhancement. Nevertheless, PM still achieves strong performance on this split, highlighting its robustness even under complex visual conditions. These results confirm that PM generalizes effectively across diverse data sources and multiple types of hallucination-oriented VQA.
@@
Performance on Shikra
To assess the generalizability of PM beyond LLaVA, we apply it, along with several comparative methods, to Shikra-7B (Chen et al. 2023), a VLM designed for referential dialogue. The performance is evaluated on MME and POPE using Shikra with iterative refinement threshold $\beta=1.5$. The results are presented in Tab. 7. Note that certain comparative methods are excluded either due to the lack of available implementations for Shikra or constraints on the computational resources available.
@@
Figure 5: Additional qualitative examples on the POPE benchmark showing the input question, the original image, the processed perception map $\mathcal{P}$, and the perception-magnified image, arranged from left to right. The answers generated by greedy decoding and PM are compared, with correct answers highlighted in green and incorrect ones in red.
Qualitative Samples of the PM
Figure 5 presents additional qualitative examples illustrating the input question, the original image, the processed perception map $\mathcal{P}$, and the perception-magnified image. The results demonstrate that the proposed method effectively enlarges small objects relevant to the question while preserving fine details. In many cases, the VLM benefits from the perception-magnified image, as it can now recognize previously overlooked objects that were too small to be detected in the original image. This leads to more accurate predictions, as observed in the first and second rows.
@@
Although region magnification allows the model to observe fine-grained visual details without altering the overall image structure, it may distort local shape properties. Specifically, in tasks requiring precise shape or length recognition, such as distinguishing letters or assessing object sizes, PM does not guarantee proportional magnification. This can introduce challenges for accurate recognition and induces hallucinations. For instance, when applied to CLEVR-style questions, as shown in Fig. 6, with the query “Please estimate how many times large is the blue sphere compared to the red cube?”, PM magnifies both objects, causing the blue sphere to appear visually larger while distorting the shape of the other two cubes. Such unintended modifications may negatively impact final predictions. Future work could explore refinements in perception map construction or post-processing techniques to achieve more structure-preserving magnification. To address the issue of distortion, several enhancements can be integrated into our framework. One potential direction involves incorporating detected bounding boxes to constrain the perception map, thereby regularizing the magnification process and preserving object shape integrity. Another approach is to group visual tokens with shared semantic representations, enabling uniform resampling across coherent regions and reducing uneven deformation within individual objects. Furthermore, for tasks where geometric fidelity is critical—such as those involving precise reasoning over shape or size—magnification can be applied adaptively at selective decoding steps rather than continuously, allowing better control over the trade-off between visual detail enhancement and structural preservation.
Efficiency.
A key limitation of our method lies in its computational overhead: applying perception-based magnification at every decoding step requires feature extraction from a modified image, which precludes the use of efficient VLM decoding techniques such as key-value (KV) caching. Nonetheless, our method remains relatively efficient when used without iterative refinement, as the perception map can be constructed with only a single additional VLM forward pass. To further mitigate runtime cost, the method can be adaptively applied only at decoding steps where visual information is critical (e.g., during object description) while being omitted for tokens with lower visual dependency, such as punctuation or reasoning-related words. A practical indicator for such adaptation could be the aggregated attention weights on visual tokens, signaling when the model is actively grounding its predictions in the visual input. Additionally, to further enhance efficiency, future work may explore speculative decoding strategies, where perception magnification is applied post hoc to refine previously decoded sequences. This deferred application may significantly reduce the frequency of image resampling, improving scalability for longer sequences without sacrificing generation quality.
@@
While the proposed method can be easily applied to interleaved VLMs with visual tokens corresponding to all patch features, applying it to VLMs where the correlation between visual tokens and image regions are non-trivial, can be slightly more complicated. The method depends on certain mapping from the visual token attention weights to the attention span on the original image. Therefore, when apply PM to models where the visual tokens are constructed with not only a feature extractor, but also a token sampler or cross-attention feature extractors, PM will need some attention mapping mechanism between visual tokens and patches. However, we can resolve this problem by introducing gradient backpropagation to address the implicit relationships between the image and tokens, refining the visual attentions and construct the perception map, similar to method proposed in (Chefer, Gur, and Wolf 2021). Overall, such limitation does not reduce the contribution of this work, as the idea of adaptive-magnification can also always be used in aligned with interpretability tools designed for different VLMs. Alternatively, one could use PM in ways similar to (Yu, Yu, and Wang 2024), where any compatible model could be used as external guidance to the decoding-adaptive magnification.
Prompt for Reference-based Evaluation
[Context]
An aerial view of Diamond Head in the Hawaiian Islands.
[Question]
What is the name of this famous sight in the photo?
[Assistant 1]
The famous sight in the photo is Diamond Head.
[End of Assistant 1]
[Assistant 2]
The famous sight in the photo is the Na Pali Coast, which is a scenic coastal area on the island of Kaua’i in Hawaii.
[End of Assistant 2]
[System]
We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above. The user asks the question on observing an image. For your reference, the visual content in the image is represented with a few sentences describing the image.
Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space.
In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgement.
Response:
10 2
Assistant 1 correctly identified the famous sight in the photo as Diamond Head, which is consistent with the context provided. The response is accurate, relevant, and directly answers the question with the appropriate level of detail given the context.
Assistant 2, on the other hand, incorrectly identified the sight as the Na Pali Coast, which is not mentioned in the context. This response is inaccurate and irrelevant to the question, as it does not align with the description of the image provided. Additionally, the details about the Na Pali Coast, while informative, are not applicable to the context of Diamond Head. Therefore, Assistant 2’s response is not helpful in this scenario.
@@
GPT-Assisted Evaluation
We adopt the GPT-assisted evaluation prompts directly from prior studies. For reference-based evaluation, we follow the prompt used in (Liu et al. 2023), with an example of the prompt and corresponding evaluation result provided in Tab. 8. For vision-based evaluation, we use the prompt from (Huang et al. 2024), with an example shown in Tab. 9.
Evaluation Protocol
Each tested method is evaluated under two settings:
  • Reference-based Evaluation: GPT-4o is given the question along with a detailed textual description of the image.
  • Vision-based Evaluation: GPT-4o is given the question and the original image.
In both settings, GPT-4o scores two answers: one from the tested method and another from GPT-4 (used as the reference). Each answer is rated independently on a scale of 0 to 10. The final evaluation score is computed as:
$\text{Relative Score}=\frac{\text{Score}_{\text{method}}}{\text{Score}_{\text{GPT-4}}}$
This normalization ensures comparability across questions of varying difficulty.
Evaluation Criteria
The evaluation criteria differ slightly between the two settings:
  • In reference-based evaluation, GPT-4o assesses overall content quality, coherence, and informativeness of the response relative to the image description.
  • In vision-based evaluation, GPT-4o is explicitly instructed to focus on the accuracy of visual grounding and the level of detail present in the answer.
Prompt for Vision-based Evaluation
You are required to score the performance of two AI assistants in describing a given image. You should pay extra attention to the hallucination, which refers to the part of descriptions that are inconsistent with the image content, such as claiming the existence of something not present in the image or describing incorrectly in terms of the counts, positions, or colors of objects in the image. Please rate the responses of the assistants on a scale of 1 to 10, where a higher score indicates better performance, according to the following criteria:
1: Accuracy: whether the response is accurate with respect to the image content. Responses with fewer hallucinations should be given higher scores.
2: Detailedness: whether the response is rich in necessary details. Note that hallucinated descriptions should not count as necessary details.
Please output the scores for each criterion, containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. Following the scores, please provide an explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgement.
[Assistant 1]
The intended effect of this painting might be to create a humorous and whimsical twist on the famous Mona Lisa artwork by featuring a dog dressed in her attire, which can evoke feelings of amusement and curiosity.
[End of Assistant 1]
[Assistant 2]
The intended effect of this painting is to create a humorous and whimsical scene by combining a dog with a human-like appearance, such as a Renaissance-style painting. The dog is wearing a Renaissance-style dress and a hat, which adds a touch of humor and creativity to the image. The painting might be intended to evoke amusement and curiosity in the viewer, as it is not common to see a dog dressed in such a manner. The artist might have wanted to challenge the viewer’s expectations and create a unique, memorable piece of art.
[End of Assistant 2]
Output format:
Accuracy: <Scores of the two answers>
Reason:
Detailedness: <Scores of the two answers>
Reason:
Response:
“9”: **Accuracy: 8 9**
*Reason:**
Assistant 1 accurately describes the image as a humorous twist on the Mona Lisa featuring a dog, but lacks specific details about the attire. Assistant 2 provides a more detailed and accurate description, mentioning the Renaissance-style dress and the humorous intent, which aligns well with the image content.
*Detailedness: 6 9**
*Reason:**
Assistant 1 gives a brief description focusing on the humorous aspect but lacks detail about the attire and style. Assistant 2 offers a richer description, noting the Renaissance-style dress and hat, and elaborates on the artistic intent, providing a more comprehensive understanding of the image.”